OpenStudy (kaylala):

proving identities: (cot+cos)/(sec-cos) = (csc + 1) / tan^2

OpenStudy (kaylala):

@Callisto @satellite73

OpenStudy (kaylala):

@jagatuba

OpenStudy (kaylala):

@dmezzullo

OpenStudy (kaylala):

@abb0t

OpenStudy (kaylala):

@Loser66 @superdavesuper

OpenStudy (kaylala):

@Preetha

OpenStudy (kaylala):

@shamil98

OpenStudy (callisto):

\[\text{L.S.}\]\[=\frac{(\cot+\cos)}{(\sec-\cos)}\]\[=\frac{\cot+\cos}{\frac{1}{\cos}-\cos}\]\[=\frac{\cot+\cos}{\frac{1-\cos^2}{\cos}}\]\[=\frac{\cot+\cos}{\sin(\tan)}\]Divide both numerator and denominator by cos\[=\frac{\frac{\cot+\cos}{\cos}}{\tan^2}\]\[=\ ...\]

OpenStudy (kaylala):

got it! thank you!! @Callisto

OpenStudy (callisto):

Welcome :)